QUESTION 11 - 1 POINT Given z = -24 - 18i, find [z]. Provide your answer below:
Answer #1

- Y -18 Given z=-24-18 it is of the form x+yi Here x=-24 Iz) - Vox?ty = V (-24)*+ (-1852 ->121= 307